OpenStudy (anonymous):

Solve the equation for y if |y + 3| = 9 y = 6, -12 y = 6, -6 y = 12, -12 y = 6

OpenStudy (anonymous):

help i need a awnser.

OpenStudy (abmon98):

\[\left| y+3 \right|=9\] \[y+3=9\] or \[y+3=-9\] \[y=6\] or\[y=-12\]

OpenStudy (anonymous):

its 12,-12

OpenStudy (anonymous):

|y+3|=9 part one when y>-3 (y+3) is always positive so y+3=9 i.e y=6 when y<-3 (y+3) is negative as absolute function (mod) is there it makes it positive -(y+3)=9 so y=-12

OpenStudy (anonymous):

its A